OpenStudy (xoxolove456):

PLEASE HELP!!! And help with steps The hypotenuse of a right triangle measures 5cm. One leg of the triangle is 3cm longer than the other. What is the length of each leg? What is the perimeter of this triangle ???

OpenStudy (pythagoras123):



OpenStudy (pythagoras123):

By Pythagoras Theorem \[x^2+(x+3)^2=25\] \[x^2 +x^2 + 9 + 6x =25\] \[2x^2+6x -16=0\] Solve for x

OpenStudy (xoxolove456):

How did you get -16?

OpenStudy (xoxolove456):

Nvm

OpenStudy (pythagoras123):

9-25=-16